OpenStudy (anonymous):

Flying to Los Angeles with a tailwind a plane averages 158km/h. On the return trip the plane only averaged 112km/h while flying back into the same wind. Find the speed of the wind and the speed of the plane in still air.

OpenStudy (anonymous):

158 = x + y 112 = x - y x is speed of plane y is speed of wind